Session: Jun 2010 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Bo Xia at Thu May 27 15:15:28 2010) Success:

Yebes (Paco Colomer at Thu May 27 16:32:26 2010) Failure: ANT
The antenna cannot be moved. This experiment has not been observed.

Onsala (Michael Lindqvist at Thu May 27 17:47:13 2010) Success:
Some light rain at the start. Otherwise no known problems.

Medicina (Roda at Thu May 27 21:10:10 2010) Success:

Urumqi (Wenjun Yang at Fri May 28 03:49:22 2010) Failure: RECO
The ftp data size is smaller than normal size, and scan_check is error. we switch bank A to bank B in ep022a exp, scan_check is ok.The vsn of n10c2 exp is med-0049, we condition that disk model now.

Torun (Kaz Borkowski at Fri May 28 07:09:46 2010) Minor failures:
We experienced a power loss at 11:30 UTC. After recovery, at 12:15, there were still fringes but about 20 times weeker than before. Reason yet unknown.

Westerbork (Geert Kuper at Fri May 28 12:46:35 2010) Success:
12 Telescopes in the array. No known problems.

Effelsberg (Uwe Bach at Wed Jun 02 07:34:06 2010) Success: